Celsius temperature readings can be converted to Fahrenheit readings using the formula F=9/5C+32. What is the Fahrenheit temperature that corresponds to each of the following Celsius temperatures: -10, 0, 15, 100?
Answer by checkley71(8403)

F=9/5C+32
F=9/5(-10)+32
F=-90/5+32
F=-18+32
F=14 ANSWER.
F=9/5(0)+32
F=0+32
F=32 ANSWER.
-------------------------------------
I'LL LEAVE THE REST FOR YOU TO SOLVE USING THE SAME PROCEDURE.